OpenStudy (anonymous):

Use basic identities to simplify the expression. cos θ - cos θ sin2θ

OpenStudy (anonymous):

sin2x=2sinxcosx

OpenStudy (anonymous):

Im sorry I still dont understand can you please explain it a little more?

OpenStudy (anonymous):

i din't get to the answer yet buy i want us to think together

OpenStudy (anonymous):

do u have any suggession?

OpenStudy (anonymous):

I need help understanding the problem in general lol so i could explain to me what you understand so far?

OpenStudy (anonymous):

okay since we have a complex problem we need to simplify it using ideentities to get a very simplified form

OpenStudy (anonymous):

Okay so what steps do we use to figure that out?

OpenStudy (anonymous):

normaly we change the complex part in the given problem to a very comprehensive one, in this cas sin2x

OpenStudy (anonymous):

Okay why is it sin2x?

OpenStudy (anonymous):

it's acctually \[\sin 2\]

OpenStudy (anonymous):

i mean \[\sin 2\theta\]

OpenStudy (anonymous):

Okay so what is the next step?